Monthly Archives: June 2014

140630 ISB

Robert arrived at ISB at 9:15. The doors opened but no sample had been collected. The top plate, drip plate, and the strand frames were removed for cleaning. The collector was disassembled and the body and collection components washed with 409 and then DI. The collecting bottle was rinsed with DI and then reattached. The system still appears to be running normally and was reset for collection of another sample.

140629 UCSC Observations

Kona arrived at ISB at 9:15 am.  The collector doors were open, no sample collected, and the system was running fine.  She took out and washed the strand frames, tray and collecting bottle (the bottle was only rinsed with DI).  The rest of the apparatus was washed with 409 and DI water.  She ran a test run and then activated the system.  Ready for fog collection.

140629 LML

Robert arrived at the LML site at 9:15. The collector had been unplugged so the doors didn’t open and no sample was collected. The top plate, drip plate, and the strand frames were removed for cleaning. The body was rinsed with DI water and the fog collector components were then cleaned with 409 followed by thorough rinsing with DI water. The system was plugged back in, keep an eye out to see if it happens again.

140628 LML

Robert arrived at the LML site at 9:30. The doors to the fog collector had opened but no sample collected. The top plate and the strand frames were removed for cleaning. The body was rinsed with DI water and the fog collector components were then cleaned with 409 followed by thorough rinsing with DI water. System was reset for collection of next sample.

140627 LML

Robert arrived at the LML site at 8:30. The doors to the fog collector had opened and less than 5 ml was in the bottle, which was discarded. The top plate and the strand frames were removed for cleaning. The body was rinsed with DI water and the fog collector components were then cleaned with 409 followed by thorough rinsing with DI water. System was reset for collection of next sample.

140625 LML Observations

Isaac arrived at the LML site at 10:30. The doors to the fog collector had opened although no visible sample had been obtained. The top plate, drip tube assembly, and the strand frames were removed for cleaning. The body and fog collector components were then sprayed initially with 409 followed by thorough rinsing with DI water. The system was then reassembled for collection during the following day.

140628-Montara

6/26/14: After acid wash redeployed Montara active sampler.

6/27/14: Sample bottle found on ground, slipped out of sample tube. System checked and rinsed and new sample bottle deployed.

6/28/14: 100 mL sample collected, preserved and refrigerated. System was still  quite clean after just one day so we just removed and cleaned the racks in DI water  and gave everything a thorough DI rinse. Collected a blank, reset system.

140625 LML Observations

Isaac arrived at the LML site at 11:10am. No sample had been collected although the doors to the fog collector had opened. The top plate, drip tube assembly, and the strand frames were removed for cleaning. The parts were initially sprayed with 409 then thoroughly rinsed with DI water. The collector body was then sprayed lightly with 409 followed by a thorough rinse with a large volume of DI water. The system was then reassembled for collection during the following day.

140625 UCSC Observations

Kona arrived at the ISB greenhouse at 9:00 am.  There was no fog sample collected and the collector doors had NOT opened; the system had crashed.  She manually restarted the system by unplugging the main power and followed with a test run which was successful.  She removed the collecting bottle and rinsed it with DI water.  The apparatus with everything else intact was spray rinsed with DI water.  The system was then activated and ready for fog collection.

140626 UCSC Observations

Jeremiah.

This morning the fog collector on the roof of ISB was open and running. There was no sample. I dismantled and cleaned the fog collector then took a blank (B-UCSC-140626). The collector is in very good condition, some spiders were found inside however it was overall clean. Log book has been returned to core lab and is in the bucket with the keys to ISB.